A company just paid a quarterly dividend of $1.00. The dividends are expected to grow by 0.5% each quarter. If the market requires an annual return of 9% on the company, what should be the company's stock price?
Answer #1

Solution

Stock price=D0*(1+g)/(r-g)

Where D0=Current dividend

g=growth rate per period=.5% per quarter

r=required rate of return per period=9/4=2.25% per quarter

Stock price=1*(1+.005)/(.0225-.005)

=57.42857
Stock price=$57.43

If you are satisfied with answer,please give a thumbs up